EXERCISE HINTS: GETTING STARTED I PM STUCK Suppose a droplet of unknown mass remains suspended against gravity when E,--3.30 x 105 Nyc What is the minimum mass of the droplet? (Take the +y direction to be upward.) The mass of an object is never negative. kg Need Help? mReedi.İ Submit Answer Save Progress Practice Another Version
Answer #1

in order to get force equilibrium to suspend droplet against gravity,

Gravitational force must balance electric force

mg = q E

m * 9.8 = q* 3.3*10^5

m = 3.367*10^4 * q

where, q = charge of droplet.

Note : as the value of charge (q) is not given so just put the value of charge (q) in the above equation to get the final value of mass (m) or just comment its value in comment section.

========

do comment in case any doubt, will reply for sure.. Goodluck